1) A company just paid a dividend of $1.50 on its stock. The dividend is expected to grow at 4% forever. If the discount rate is 6%, what is the present value of the stock?

Group of answer choices

$80.97

$74.00

$79.38

$78.00

2)

A stock is expected to pay a dividend of $3 next year. The dividend will grow at a rate of 5% for 2 years, and will then grow at a rate of 3% from that point on. If the required rate of return is 7%, what is the stock’s value today?

Group of answer choices

$77.78

$79.03

$84.01

$75.55

3)A company issues preferred stock that pays a dividend of $3.25 indefinitely. If an investor's required return is 7%, what should the price of the preferred stock be?

Group of answer choices

$49.22

$45.98

$46.43

$47.50

4)

A company will pay a dividend of $2 per share next year. If the current price of the stock is $39 and the growth rate of the dividend is 3%, what is the required return on this stock?

Group of answer choices

8.68%

9.03%

8.13%

7.56%
Answer #1

1 D1 DO*(1+g) = 1.50*(1+0.04) = 1.56 Share Price Next Year dividend/(Cost of Equity-Growth rate) 1.56/(0.06-0.04) $ 78.00

3.00% 7.00% Pv Factor Present Value 0.9346 $ 2.80 0.8734 $ 2.75 0.8163 $ 2.70 $ 3.00000 $ 3.15000 $ 3.30750 $ 3.40673 $ 85.16

Preferred Stock 3 Price Annual Dividend/Required Return 3.25/0.07 $ 46.43 4 Required return (Next Year Dividend/Price) +Growt